OpenStudy (anonymous):

Marvel Man movie is being released to a limited number of movie theaters. The number of showtimes for the movie is twice the number of movie theaters. The number of seats available for each showtime is three times the number of showtimes. Let x represent the number of theaters showing the movie.

OpenStudy (anonymous):

Could it be set up like this. x^2+s^3

OpenStudy (anonymous):

boooooooooooooooooo

OpenStudy (anonymous):

flop movie

OpenStudy (anonymous):

collected $0 at the box office

OpenStudy (anonymous):

ranga is typing a reply…

OpenStudy (anonymous):

wtf!

OpenStudy (ranga):

# of theaters = x # of showtimes = 2*x (because number of showtimes for the movie is twice the number of movie theaters) # of seats = 3 * 2 * x (because number of seats available for each showtime is three times the number of showtimes)

OpenStudy (anonymous):

3*2*x would be the expression for that?

OpenStudy (ranga):

Yes, the number of seats = 3*2*x = 6x.

OpenStudy (ranga):

But the question seems incomplete. They are not asking for anything.

OpenStudy (anonymous):

Well yea

OpenStudy (anonymous):

A. Determine the three factors that are to be multiplied to determine the total number of seats available for the movie. B. When simplified what is the degree and classification of this expression? C. Interpet each pairing of factors as a single unit in the context of the problem I think I could manage a and b but how would I get C?

OpenStudy (ranga):

A) The three factors to be multiplied are x, 2 and 3 B) when a) is simplified we get 6x. This is polynomial of degree 1 (because the highest exponent of x is 1). It is also a monomial because there is only one term C) The factors are: x, 2 and 3. If we pair x and 2, the product is the number of showtimes. You can see if other pairings can be similarly interpreted or not.

OpenStudy (anonymous):

Ah thanks man

OpenStudy (anonymous):

Big help